# Towards Imperceptible JPEG Deep Hiding: Multi-range Representations-driven Adversarial Stego Generation

### v1

@@ 3.3 Loss Function
The total loss consists of three parts: the hiding loss to ensure visual imperceptibility, the revealing loss to ensure secret restorability, and the features angle-norm disentanglement loss to guarantee steganalysis imperceptibility. is equal to the sum of these three losses, expressed as follows
In Eq. (4), represents the actual restored secret image.
Method | COCO | ImageNet | BOSSBase | ||||||
ACC_U | ACC_E | ACC_D | ACC_U | ACC_E | ACC_D | ACC_U | ACC_E | ACC_D | |
EFDR Yang and Liao (2023) | 92.73 | 75.68 | 99.90 | 92.85 | 75.65 | 99.90 | 87.65 | 73.70 | 99.98 |
MRAG (UCNet) | 50.05 | 51.28 | 74.53 | 50.00 | 51.68 | 78.83 | 50.00 | 50.55 | 63.28 |
MRAG (EWNet) | 58.53 | 60.48 | 71.60 | 60.73 | 59.25 | 74.65 | 51.20 | 55.00 | 59.05 |
MRAG (DBS2Net) | 53.35 | 51.88 | 50.13 | 53.50 | 51.18 | 50.30 | 51.25 | 50.43 | 50.13 |
(a) UCNet_EFDR
(c) EWNet_EFDR
(e) DBS2Net_EFDR
(b) UCNet_MRAG
(d) EWNet_MRAG
(f) DBS2Net_MRAG
The Number of Layer | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | #Param () | ||||
PSNR | SSIM | PSNR | SSIM | ACC_U | ACC_E | ACC_D | ||
1 | 41.59 | .987 | 33.21 | .941 | 50.05 | 52.48 | 75.88 | 14.54 |
2 | 40.56 | .984 | 32.60 | .934 | 50.08 | 52.85 | 98.90 | 14.73 |
3 | 40.11 | .983 | 31.68 | .920 | 50.05 | 51.28 | 74.53 | 14.91 |
4 | 39.72 | .983 | 31.80 | .923 | 50.05 | 64.23 | 80.30 | 15.10 |
Local Range | Global Range | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | ||||
PSNR | SSIM | PSNR | SSIM | ACC_U | ACC_E | ACC_D | ||
✗ | ✔ | 26.23 | .786 | 31.95 | .924 | 50.03 | 51.15 | 67.75 |
✔ | ✗ | 36.20 | .965 | 31.27 | .908 | 50.18 | 51.10 | 85.13 |
✔ | ✔ | 40.11 | .983 | 31.68 | .920 | 50.05 | 51.28 | 74.53 |
4 Experiment
4.1 Experimental Settings
@@
We employ three image datasets in our implementation, including COCO Lin et al. (2014), ImageNet Russakovsky et al. (2015), which are widely used in various deep learning tasks, and a standard steganographic dataset BOSSBase Bas et al. (2011). We utilize 5000, 1000, and 2000 cover-secret pairs from COCO for training, validation, and testing, respectively. 2000 cover-secret pairs from ImageNet and BOSSBase, respectively, are used only for testing. Additional 10000 cover-stego pairs, randomly mixed from three datasets, are used for training the steganalyzer. The stego images are generated by HRJS Yang et al. (2023), a bit-level JPEG steganographic method based on deep learning. All images are center-cropped to . The compression quality factor adopted by the cover is 75, the most commonly used in JPEG steganography.
@@
We take several advanced and representative methods as our baselines, including ISN Lu et al. (2021a), a DH-Spatial method, RIIS Xu et al. (2022), a robust DH-Spatial method, and EFDR Yang and Liao (2023), which is currently the only customized deep hiding method with color JPEG images as covers. Note that during the training phase of ISN, we add an additional JPEG compression distortion layer, referred to as ISN+ for distinction.
@@
To measure the secret restorability and visual imperceptibility, we adopt peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) Wang et al. (2004), average pixel discrepancy (APD), and learned perceptual image patch similarity (LPIPS) Zhang et al. (2018a). PSNR/SSIM with larger values and APD/LPIPS with smaller values indicate higher image quality. For steganalysis imperceptibility, we utilize the detection accuracy rate (ACC) as the metric. The ACC is closer to 50%, i.e., random guess, indicating superior performance. ACC_U, ACC_E, and ACC_D denote the detection accuracy rate of UCNet, EWNet, and DBS2Net, respectively.
@@
Before training MRAG, the surrogate steganalyzer is pre-trained 50 epochs on 5000 mixed cover-stego pairs with the parameters following the default settings. Subsequently, the pre-trained surrogate steganalyzer’s weights are frozen, and we train MRAG for 200 epochs on 5000 cover-secret pairs from COCO. We adopt the Adam optimizer with , , and , and set the learning rate to . The mini-batch size is 4, containing two randomly selected cover-secret pairs. Note that the steganalyzer used for the steganalysis imperceptibility evaluation is trained on another 5000 mixed cover-stego pairs. All experiments are implemented with PyTorch and on the NVIDIA GeForce RTX 4090 GPU.
4.2 Comparison Results
In this subsection, to verify the performance of MRAG in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility, we conduct extensive experiments on three datasets and three open-source JPEG steganalyzers.
@@ 4.2 Comparison Results
Table 1 reports the numerical comparisons of our MRAG with existing baselines. It can be observed that MRAG significantly outperforms these baselines in terms of secret restorability. Compared with EFDR, MRAG can supply at least 2.95 dB PSNR advancement. Similar performance improvements can also be observed in SSIM, APD, and LPIPS. Visualization comparisons are shown in Fig. 3. We can observe that the secret images revealed by MRAG are visually pleasing with minimal visual artifacts. In contrast, the DH-Spatial methods ISN+ and RIIS yield visible color distortion, blocking, blurring, and banding artifacts in the revealed secret images. It is noteworthy that although our model only uses the COCO for training, it still provides excellent results on the other two datasets, which indicates that our model has a favorable generalization ability.
@@
Table 2 and Fig. 4 report the numerical comparisons and visualization comparisons of our MRAG with ISN+, RIIS, and EFDR in terms of visual imperceptibility. We can observe that MRAG and EFDR have comparable performance, while the stego images generated by ISN+ and RIIS still contain many visible artifacts, making it difficult for them to achieve visual imperceptibility.
@@
Steganalysis imperceptibility, an essential metric, evaluates the statistical imperceptible to resist steganalyzers. It achieves classification by distinguishing the changes in statistical features caused by hiding. Table 3 shows the numerical comparisons of our MRAG with EFDR. It can be observed that MRAG significantly outperforms EFDR in terms of steganalysis imperceptibility. Compared to EFDR, MRAG can provide at least 15.2% improvement, with most classification results approaching 50%. Moreover, we visualize the distribution of the classified features learned by three steganalyzers using t-distributed stochastic neighbor embedding (t-SNE) in Fig. 5. We can find that the distribution of covers’ classified features (green points) and stegos’ classified features (red points) obtained by EFDR can be well distinguished, while the distribution of the classified features obtained by the proposed MRAG is intermingled and difficult to distinguish. This also confirms the superior steganalysis imperceptibility of MRAG.
4.3 Ablation Study
Here, we mainly discuss the impact of the number of the affine coupling layer and multi-range branches.
@@ 4.3 Ablation Study
Also, we verify the performance effects of the invertible local-range branch and the invertible global-range branch, as shown in Table 5. It can be observed that compared to any single-range branch, our designed multi-range branches can significantly boost the performance of secret restorability and steganalysis imperceptibility while keeping visual imperceptibility.
5 Conclusion
In this work, we attempt to introduce generation-based adversarial attacks into color JPEG image deep hiding, exploring the potential of generated color JPEG stego images in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility. To the best of our knowledge, this is the first attempt to focus on the steganalysis imperceptibility issue of color JPEG image deep hiding. A multi-range representations-driven adversarial stego generation framework, MRAG, is proposed from a steganalysis perspective. Besides, a features angle-norm disentanglement loss is designed to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features. Comprehensive experiments verify the effectiveness of MRAG and its superior performance. We believe the success of MRAG can further promote the development of color JPEG image deep hiding and provide some inspiration for further research on steganalysis imperceptibility.

### v2

@@ Introduction
The total loss consists of three parts: the hiding loss to ensure visual imperceptibility, the revealing loss to ensure secret restorability, and the features angle-norm disentanglement loss to guarantee steganalysis imperceptibility. is the sum of these three losses, expressed as follows
where , and . represents the actual restored secret image.
| Method | COCO | ImageNet | BOSSBase | |||||||||
| A_U | A_E | A_D | Average | A_U | A_E | A_D | Average | A_U | A_E | A_D | Average | |
| EFDR (Yang and Liao 2023) | 92.73 | 75.68 | 99.90 | 89.44 | 92.85 | 75.65 | 99.90 | 89.47 | 87.65 | 73.70 | 99.98 | 87.11 |
[…]
Experiments
Experimental Settings
@@
We employ three image datasets in our implementation, including COCO (Lin et al. 2014), ImageNet (Russakovsky et al. 2015), which are widely used in various deep learning tasks, and a standard steganographic dataset, BOSSBase (Bas, Filler, and Pevný 2011). We utilize 5000, 1000, and 2000 cover-secret pairs from COCO for training, validation, and testing, respectively. 2000 cover-secret pairs, each from ImageNet and BOSSBase, are used only for testing. Additional 10000 cover-stego pairs, randomly selected from three datasets, are used to train the steganalyzer. The stego images are generated by HRJS (Yang et al. 2023), a bit-level JPEG steganographic method based on deep learning. All images are center-cropped to and , due to the limit of computational resources. The compression quality factor adopted by the cover is 75, the most commonly used in JPEG steganography. The code will be open-sourced on GitHub upon acceptance.
@@
We use several advanced and representative methods as our baselines, including ISN (Lu et al. 2021), a SIH method, RIIS (Xu et al. 2022), a robust SIH method, and EFDR (Yang and Liao 2023) that is currently the only customized image hiding method with color JPEG images as covers. Note that during the training phase of ISN, we add an additional JPEG compression distortion layer, referred to as ISN+ for distinction.
@@
To measure the secret restorability and visual imperceptibility, we adopt peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) (Wang et al. 2004), average pixel discrepancy (APD), and learned perceptual image patch similarity (LPIPS) (Zhang et al. 2018). Higher PSNR/SSIM values and lower APD/LPIPS indicate higher image quality. For steganalysis imperceptibility, we utilize the detection accuracy rate as the metric. The detection accuracy rate is closer to 50%, i.e., a random guess, indicating superior performance. A_U, A_E, and A_D denote the detection accuracy rate of UCNet, EWNet, and DBS2Net, respectively.
@@
Before training MRAG, the surrogate steganalyzer is pre-trained for 50 epochs on 5000 mixed cover-stego pairs following default settings. Subsequently, the pre-trained surrogate steganalyzer’s weights are frozen, and we train MRAG for 200 epochs on 5000 cover-secret pairs from COCO. Note that the steganalyzer used for the steganalysis imperceptibility evaluation is trained on another 5000 mixed cover-stego pairs. We adopt the Adam optimizer with , , and , and set the learning rate to . The mini-batch size is 4, containing two randomly selected cover-secret pairs. All experiments are implemented with PyTorch on the NVIDIA GeForce RTX 4090 GPU.
| The Number of Layer | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | |||||
| PSNR | SSIM | PSNR | SSIM | A_U | A_E | A_D | Average | |
| 1 | 41.59 | .987 | 33.21 | .941 | 50.05 | 52.48 | 75.88 | 59.47 |
| 2 | 40.56 | .984 | 32.60 | .934 | 50.08 | 52.85 | 98.90 | 67.28 |
[…]
| 4 | 39.72 | .983 | 31.80 | .923 | 50.05 | 64.23 | 80.30 | 64.86 |
| Local | Global | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | |||||
| PSNR | SSIM | PSNR | SSIM | A_U | A_E | A_D | Average | ||
| ✗ | ✔ | 26.23 | .786 | 31.95 | .924 | 50.03 | 51.15 | 67.75 | 56.31 |
| ✔ | ✗ | 36.20 | .965 | 31.27 | .908 | 50.18 | 51.10 | 85.13 | 62.14 |
[…]
Comparison Results
To verify the performance of MRAG in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility, we conduct extensive experiments on three datasets and three open-source JPEG steganalyzers.
@@
Table 1 reports the numerical comparisons of MRAG and baselines. It can be observed that MRAG significantly outperforms these baselines in terms of secret restorability. Compared with EFDR, MRAG can supply at least 2.95 dB PSNR advancement. Similar performance improvements can also be observed in SSIM, APD, and LPIPS. Visualization comparisons are shown in Fig. 3. We can observe that the secret images revealed by MRAG are visually pleasing with minimal visual artifacts. In contrast, the SIH methods ISN+ and RIIS yield visible color distortion, blocking, blurring, and banding artifacts in the revealed secret images. It is noteworthy that although our model only uses the COCO for training, it still provides excellent results on the other two datasets, which indicates that our model has a favorable generalization ability.
@@
Table 2 and Fig. 4 report the numerical comparisons and visualization comparisons of our MRAG with ISN+, RIIS, and EFDR in terms of visual imperceptibility. We can observe that MRAG and EFDR have comparable performance, while the stegos generated by ISN+ and RIIS still contain many visible artifacts, making it difficult for them to achieve visual imperceptibility.
@@
Steganalysis imperceptibility is a more rigorous imperceptibility metric, evaluating the ability to resist steganalyzers. It achieves classification by distinguishing the changes in statistical features caused by hiding. Table 3 shows the numerical comparisons of our MRAG and EFDR. We can observe that MRAG significantly outperforms EFDR. Compared to EFDR, MRAG can provide at least 15.2% improvement, with most classification results approaching 50%. Moreover, we visualize the distribution of the classified features learned by three steganalyzers using t-distributed stochastic neighbor embedding (t-SNE) in Fig. 5. We can find that the distribution of covers’ classified features (green points) and stegos’ classified features (red points) obtained by EFDR can be well distinguished, while the distribution of the classified features obtained by the proposed MRAG is intermingled and difficult to distinguish. This also confirms the superior steganalysis imperceptibility of MRAG.
Due to the page limit, comparison results with a size of can be found in the supplementary materials.
Ablation Study
Here, we mainly discuss the impact of the number of the affine coupling layer and multi-range branches. The image size used in ablation is , unless otherwise specified.
@@
Also, we verify the performance effects of the invertible local-range branch and the invertible global-range branch, as shown in Table 5. It can be observed that compared to any single-range branch, the designed multi-range branches can significantly boost the performance of secret restorability and steganalysis imperceptibility while keeping visual imperceptibility.
More experiments include ablation of features angle-norm disentanglement loss and conventional cross-entropy loss, ablation of coarse/fine-grained maps feeding into local/global branches, and comparisons of computational overhead can be found in the supplementary materials.
Conclusion
This work is the first attempt to explore the potential of the generated color JPEG stego in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility. It designs a novel adversarial stego generation framework, MRAG, from a steganalysis perspective. Besides, a features angle-norm disentanglement loss, based on the surrogate steganalyzer’s classified features, is incorporated to launch multi-range representations-driven feature-level adversarial attacks. It constrains MRAG to delicately encode the concatenation of cover and secret into subtle adversarial perturbations from local and global ranges relevant to steganalysis. Comprehensive experiments verify the effectiveness and superiority of MRAG. We believe the success of MRAG can further promote the development of color JPEG image hiding and provide some inspiration for further research on steganalysis imperceptibility.